### v1

# Vamos: Versatile Action Models for Video Understanding

Abstract
What makes good video representations for video understanding, such as anticipating future activities, or answering video-conditioned questions? While earlier approaches focus on end-to-end learning directly from video pixels, we propose to revisit text-based representations, such as discrete action labels, or free-form video captions, which are interpretable and can be directly consumed by large language models (LLMs). Intuitively, different video understanding tasks may require representations that are complementary and at different granularities. To this end, we propose versatile action models (Vamos), a learning framework powered by a large language model as the “reasoner”, and can flexibly leverage visual embeddings, action labels, and free-form descriptions extracted from videos as its input. We evaluate Vamos on four complementary video understanding benchmarks, Ego4D, Next-QA, IntentQA, and EgoSchema, on its capability to model temporal dynamics, encode visual history, and perform reasoning. Surprisingly, we observe that text-based representations consistently achieve competitive performance on all benchmarks, and that visual embeddings provide marginal or no performance improvement, demonstrating the effectiveness of text-based video representation in the LLM era. We perform extensive ablation study and qualitative analysis to support our observations, and achieve state-of-the-art performance on three benchmarks. Code and models will be released at https://brown-palm.github.io/Vamos/.
1 Introduction
Figure 1: Visual observations with vastly diverse appearances may be described with the same captions. Our work explores the use of general-purpose text descriptions of video data for action anticipation and question answering. We propose Vamos, a versatile reasoning framework that allows us to study the impact of latent visual representations, discrete action labels, and free-form text descriptions for downstream applications.
A picture held us captive … for it lay in our language and language seemed to repeat it to us inexorably.
Philosophical Investigations
Building a generative model for everyday human activities has long been a dream for researchers working on video understanding. Central to this problem are capturing the interactions between humans and the environment [67, 12], modeling the temporal dynamics of activities [65, 15], and encoding the hierarchical structures among atomic actions [59, 23], activities [6, 7], and events [37, 35]. Once constructed, the generative model of actions can be applied to a wide range of tasks, including activity and event recognition [56], future behavior prediction [49], goal and intent inference [57], and temporal reasoning [71].
Despite its desirable properties, generative modeling of actions from video observations remains challenging, hindered by two open research questions: First, what makes good video representations? Earlier attempts often relied on manually defining the actions and the objects being interacted with [23, 1, 59]. They require task-specific prior knowledge, and cannot generalize to the “open vocabulary” scenarios in the wild. Alternative approaches aim to model the temporal dynamics of human pose [51, 45, 36] or latent representations encoded by deep neural networks [65, 32], which are either too fine-grained, or not directly intepretable. Second, what makes a good model of human actions? While earlier approaches attempted to apply rule-based generative action grammars [30, 25, 54, 56], they may not be able to capture the diverse, even peculiar ways of how events would unfold over time. More recent approaches adopt a data-driven framework and directly learn autoregressive models [40] on visual tokens [62, 74], where the visual domain is often specialized (e.g. cooking, or robotics).
We aim to address both challenges by exploring an unconventional idea: Can task-agnostic natural language descriptions, such as those generated by off-the-shelf image caption models [41, 48] on sampled video frames, serve as useful video representations for action modeling from videos? And if so, can we then leverage a pre-trained large language model (LLM) [64] as the generative model of actions, represented as free-form text? Despite the apparent crudeness, we argue that it is inspired and supported by recent advances in LLM research, where LLMs have been shown to be capable of learning context-free grammar [3] with long-range dependencies, predicting time-series [22, 52], and performing reasoning [4, 13, 70, 14], all of which are indispensable for action modeling.
To rigorously validate this idea, we propose versatile action models (Vamos), a framework that flexibly unifies three representations, namely distributed visual embeddings, discrete action labels, and free-form text descriptions, and can be applied to various applications by leveraging large language models, such as Llama-2 [64]. The visual embeddings are linearly projected into the same language space following standard practice [41, 53]. As illustrated in Figure 1, the same caption can sometimes be used to describe visually diverse inputs, it is thus essential for Vamos to be able to leverage one or multiple representations simultaneously, to understand the impact of individual representation type. Vamos can directly leverage an LLM’s next token prediction capability for action anticipation [21]. Similar to how an LLM can be tasked to perform reasoning, we ask Vamos to perform video question answering [71], by prepending the question to the video representation.
We perform extensive evaluations on four benchmarks, including the Ego4D [21] long-term action anticipation, NeXT-QA [71] and IntentQA [43] video question answering, and EgoSchema [50] zero-shot long-form video question answering. We observe that for the direct application of Vamos in the action anticipation task, the representation based on discrete action labels outperforms its counterpart based on visual embeddings. We further observe that free-form video descriptors serve as an effective long-video representation that generalizes well in zero-shot setting, outperforming the strongest video-language model [68] by 50%. We then confirm that our observations are general, that text-based representation consistently provides competitive performance across all tasks, and that adding visual embeddings only result in marginal performance gains.
Another benefit of text-based video representation is interpretability. We exploit this property in two scenarios: First, we perform causal intervention and correct the suboptimal video descriptions, and we observe that the answers are indeed corrected accordingly. Second, we propose a lightweight token selector module to sample as few as 20 tokens from video descriptions that are more than 600 tokens long. We observe that not only the video question answering performance remains comeptitive, outperforming the visual embedding baseline, but the selected tokens are indeed highly related to answering the questions, confirming that the Vamos framework behaves as intended.
To summarize, we make the following contributions:
  1. 1.
    We propose to revisit video representation and action modeling in the LLM era, and to explore the effectiveness of free-form text descriptions.
  2. 2.
    We propose Vamos, a versatile framework that allows us to incorporate and compare the effectiveness of different video representations, both visual and textual.
  3. 3.
    We demonstrate the effectiveness of free-form text as a representation that is performant, interpretable, and can be intervened, even when aggressively compressed to 6%. Vamos achieves state-of-the-art performance on Ego4D LTA, NeXT-QA, IntentQA, and EgoSchema.
2 Related Work
Figure 2: Overview of Vamos and token selector. (a) Vamos flexibly unifies distributed visual features and textual video representations including discrete action labels and free-form video captions. By performing sequence modeling, Vamos can be applied to multiple video understanding tasks. (b) The lightweight token selector as an add-on module of Vamos. It pick one single token from the video sequence based on the task sequence. By applying the token selector to $k$ segments of textual video representation, $k$ tokens are finally selected as the compact input to the LLM for down-stream tasks.
Vision-Language Foundation Models. Vision-Language Models (VLMs) represent a pivotal convergence in artificial intelligence, aiming to bridge the gap between visual understanding and natural language comprehension. Recent advancements in vision-language models have been propelled by various pretraining techniques leveraging large-scale multimodal datasets. Models such as CLIP [58], ALIGN [33], CLOOB [19] bridge the vision and language modalities by learning a text encoder and an image encoder jointly with a contrastive loss, using large datasets consisting of image and caption pairs. Another line of vision-language models showcased by VisualBERT [44], UNITER [11], OSCAR [46], FLAVA [61], uses a combination of Masked-Language Modeling (MLM) and Image-Text Matching (ITM) objectives to align specific parts of images with text and enable various downstream tasks such as visual question answering, visual commonsense reasoning, text-based image retrieval, and text-guided object detection. For videos, VIOLET [17] trains an end-to-end transformer for video and language modalities, by representing videos as visual tokens and performing joint masked token modeling. To perform visual-language joint training, speech transcripts are often used as the language modality for videos [62, 17, 81, 82]. The objectives can be combined [78] and the encoders for different modalities can be shared [66]. Compared to existing vision-language models, Vamos imposes an “information bottleneck” when text-based representation is used: It converts visual inputs into discrete action labels and free-form text descriptions.
Visually-augmented LLMs. Apart from joint visual-language pre-training, existing large language models (LLMs) can also be augmented to incorporate visual inputs. For example, VisualGPT [9] and Flamingo [2] directly fuse visual information into the layers of a language model decoder using a cross-attention mechanism instead of using images as additional prefixes to the language model. Other approaches, such as instructional tuning [48], or linearly projecting the visual embeddings into the input space of LLMs [41, 53], have also been explored. Vamos largely follows this linear projection approach to incorporate visual embedding inputs, with the goal to understand if and how they are complementary to text-based video representations.
Additionally, tool-using large language models have been recently proposed to invoke and incorporate the use of task-specific modules [60, 27], where visual perceptions consist a substantial subset of the tools. Notably, VisualProgram [24] and ViperGPT [63] propose to apply LLMs to generate symbolic programs based on pre-selected computer vision modules for visual question answering. Closest to our work is Socratic Models [83], where the authors propose to use natural language as the common interface to connect foundation models with different input and output modalities. Vamos differs from these approaches as it aims to understand the power and limitations of text-based representation for a wide range of video understanding tasks, through quantitative experiments and qualitative analysis.
3 Method
We now describe how the text-based representation is constructed and incorporated into versatile action models.
3.1 Text-based Video Representation
Compared with a single image, the video data often contains more complex and dynamic information including context and interactions. Prior works [39, 80, 69] have demonstrated the effectiveness of condensing images into conceptual or natural language bottleneck with good interpretability, while it is still unclear if long form videos with richer dynamics can also be condensed into such explainable textual representations. To translate videos into language bottleneck, we consider two methods: discrete action labels and free-form text descriptions.
We propose to use an extremely compact representation by condensing videos into sequences of discrete action labels, offering a more compact representation. In practice, this is achieved through the application of action recognition models such as Transformer encoders that operate in the pre-defined action space.
Video text descriptions use free-form natural language to characterize objects, scenes, and actions, which succinctly summarize the essential elements depicted in the video. It transforms complex and dynamic visual information into more comprehensible and interpretable textual descriptions. We employ off-the-shelf image captioning models such as BLIP-2 [42] that process sampled frames to produce frame-level captions, which are subsequently concatenated to form a comprehensive video-level caption.
3.2 Versatile Action Models
Benefiting from the encoded extensive prior knowledge, large language models (LLMs) have demonstrated strong capability for temporal and causal reasoning, which are crucial for video understanding. We introduce Vamos, a simple yet effective framework to utilize LLMs to unify video dynamic modeling tasks, including comprehending historical content (video question answering, VQA) and future prediction (long-term action anticipation, LTA). As shown in Figure 2 (a), given a video $V$ and a pretrained LLM $\pi$, the input sequence $x_{t}=[x_{tvr},x_{task}]$ consists of the textual video representations $x_{tvr}$ of $V$ and other task specific langauge inputs $x_{task}$ (e.g., instructions, questions, targets). The frozen word embedding layer $\mathcal{F}_{emb}$ first generate the corresponding text tokens $z_{t}=\mathcal{F}_{emb}(x_{t})\in\mathbb{R}^{L_{t}\times D}$, where $L_{t}$ is the sequence length of $x_{t}$, $D$ is the feature dimension.
To address any residual information not entirely captured by the text-based representations, we incorporate a learnable linear projection layer $\mathcal{F}_{proj}$ to align visual features with the language space. Specifically, the frozen vision backbone $\mathcal{E}$ takes in $N_{v}$ frames $[v_{1},...v_{N_{v}}]$ sampled from $V$ to generate the visual features. These visual features are then input into the projection layer $\mathcal{F}_{proj}$ to produce vision tokens $z_{v}=\mathcal{F}_{proj}(\mathcal{E}(v_{1},...v_{N_{v}}))\in\mathbb{R}^{N_{v}\times D}$. $z_{v}$ and $z_{t}$ are then concatenated and fed into the LLM $\pi$. During training, the LLM $\pi$ is optimized with the standard language modeling loss using LoRA [26].
Vamos is accommodating multiple video understanding tasks that can be formatted as sequence modeling tasks with varying task-specific inputs $x_{task}$ and modeling objectives. Specifically, for the VQA task, $x_{task}$ is composed of instructions, questions, and answers, with the answer being the training objective. During inference, the answer that maximizes sequence modeling likelihoods is selected for multiple-choice QA, or directly generated for open-ended QA, depending on the dataset. For the LTA task, $x_{task}$ is composed of instructions and future actions, where the training objective is the future action sequence. During inference, the trained LLM generates future action sequences based on history actions.
3.3 Lightweight Token Selector
We also investigate the potential of compressing long input sequences and extracting keys element from the generated textual representations to probe the crucial information for downstream video understanding tasks. We design a lightweight token selector as a add-on module (shown in Figure 2 (b)) for Vamos. The token selector takes in a sequence of textual video tokens to be selected, and a task-specific token sequence, functioning as queries. It picks a single token from the video tokens that is most relevant for downstream video understanding. To select a condensed token sequence for a video, the original textual video token sequence with $L$ tokens is first uniformly divided into $k\ll L$ segments $\{X^{(1)}$,…, $X^{(k)}\}$ containing $n$ tokens. Each segment $X^{(i)}=\{x_{1}^{(i)},...,x_{n}^{(i)}\}$ is fed into the token selector as candidate tokens, from which one token $x^{(i)}$ is selected based on the task texts (question for VQA).
Within the token selector, $\{x_{1}^{(i)},...,x_{n}^{(i)}\}$ are first projected to a lower dimension, then input into a shallow transformer encoder to obtain encodings $\{s_{1}^{(i)},...,s_{n}^{(i)}\}$. A linear layer then takes these encodings and generates the logits $g^{(i)}\in\mathbb{R}^{n}$ for final selection. During training, to maintain model differentiability, we apply Gumbel-Softmax [31] to the logits $g^{(i)}$ to pick the final selected token $x^{(i)}$ for each segment $X^{(i)}$. In this way, $k$ tokens $\{x^{(1)},...,x^{(k)}\}$ are sampled as a condensed representation of the original text sequence with much smaller length and input to downstream LLMs for video understanding.
4 Experiments
Table 1: Vamos with various video representations on LTA task. Results are reported on Ego4D LTA test set, the metric is edit distance.
Input Ego4D-LTA normal-$\downarrow$
Verb Noun Action
vision 0.653 0.673 0.884
action 0.661 0.651 0.878
caption 0.676 0.675 0.890
vis+act 0.643 0.650 0.868
Input NeXT-QA normal-$\uparrow$ IntentQA normal-$\uparrow$
Cau. Tem. Des. All CW CH TP&TN All
vision 61.5 58.2 66.3 61.2 55.2 62.4 46.9 54.4
action 55.1 53.4 53.2 54.3 57.5 65.2 47.4 56.3
caption 70.9 69.3 77.7 71.5 69.7 68.8 61.7 67.6
vis+cap 70.9 67.2 77.7 70.8 69.0 69.9 63.2 67.7
Table 1: Vamos with various video representations on LTA task. Results are reported on Ego4D LTA test set, the metric is edit distance.
Table 2: Vamos with various video representations on VQA task. Results are reported on NeXT-QA and IntentQA test set, the metric is accuracy. On both dataset, caption-based representations achieve competitive performance.
In this section, we conduct experiments on two tasks and four datasets with both quantitative and qualitative analysis.
4.1 Task and Datasets
We focus on two major video understanding tasks: long-term action anticipation and video question answering. Long-term action anticipation. The long-term action anticipation (LTA) task asks a model to predict a sequence of actions in form of a verb-noun pairs in a long future window based on video observations of the past actions. In LTA, a long video $V$ is first split into a number of annotated video segments. Given video observation before segment $i$, our task is to predict the future actions in sequences of verb-noun pairs of the next $Z$ segments allowing $K$ candidate sequences. The correctness of the predicted sequence is measured with edit distance. We evaluation on:
Ego4D [21] is comprised of 3,670 hours of egocentric videos in hundreds of scenarios of daily life activity. The Ego4D LTA v2 benchmark we focus on includes a total duration of around 243 hours of videos annotated into 3472 clips with 117 verbs and 521 nouns. We follow the official dataset splits and adopt the official parameters of the evaluation metric, with $Z=20$ and $K=5$.
Video question answering. Given a set of videos $V$, and a corresponding set of language-based questions $Q_{v}$ and their candidate answers $A_{q}$. The goal of video question answering (VQA) task is to predict the correct answer $A$ for each video-question pair. The performance is measured by accuracy. For VQA, we evaluate on three datasets:
EgoSchema [50] is a recently proposed diagnostic datasets derived from Ego4D for long-form video understanding. In contrast to previous video question answering benchmarks where the input videos are generally less than one-minute long, Egoschema contains videos that span 3-minute long with high temporal certificate to testify the video understanding model’s capability to understand long videos. It contains 5031 videos and each video has a multiple choice question. It requires zero-shot evaluation and releases answers to 500 questions for development.
NeXT-QA [71] is a popular multiple choice video question answering benchmark that test video understanding in terms of describing and reasoning the temporal actions. It contains 5,440 video clips and 47,692 questions, grouped into causal (48%), temporal (29%) and descriptive (23%).
IntentQA [43] is another multiple choice VQA dataset built on top of NeXT-QA but focuses on intent reasoning. The authors select the videos related to causal and temporal questions from NeXT-QA and constructed their own questions and answers to focus on testifying models’s performance on reasoning questions.
4.2 Implementation
Generating Action Labels. To generate action labels for videos, we use a recognition model pretrained on Ego4D LTA. It is a 3-block 6-head transformer-encoder that takes 4 CLIP features and outputs two logits for verb and noun respectively. It predicts actions in the action space pre-defined by Ego4D LTA. For each video, our action recognition model samples 4 frames for each 8s segment splits uniformly and output a verb and noun pair for the segment.
Generating Video Captions. We zero-shot generate task-agnostic video captions using BLIP-2 for Ego4D LTA, IntentQA, and EgoSchema, LLaVA-1.5 for NeXT-QA. For Ego4D LTA, we sample the center frame for each video segment to generate its caption. The captions for the 8 observed segments are then concatenated as the video representations. For VQA benchmarks, we first uniformly sample a fixed number of frames for each video, then caption each frame and concatenate these frame-level captions to represent the video. For NeXT-QA, IntentQA, and EgoSchema we sample 6, 6, and 12 frames respectively.
Vamos for Temporal Modeling. For full-shot VQA and LTA, We use Llama2-7B [64], an open-sourced LLM as the temporal model for video understanding. During training, we use low-rank adaption [26] (LoRA) adapter to perform parameter-efficient fine-tuning on the training set. For vision input, we use the frozen CLIP [58] ViT-L/14 to extract image features. For the zero-shot long-form VQA on EgoSchema, we use several popular LLMs including OpenAI GPT-3.5-turbo, GPT-4 and Llama2-chat-13B.
4.3 LTA: Action Modeling with Vamos
We first apply Vamos on the long-term action anticipation task, which requires direct modeling of video temporal dynamics by predicting future actions tokens based on video observation. We finetune Vamos on Ego4D LTA dataset with three types of representations: the recognized action labels, free-formed captions, and the continuous vision features. As detailed in the first three rows in Table 2, we observe that the action-based representation typically outperforms the vision-based input. Notably, the action representation, despite its compact nature, surpasses captions, which provide more extensive details and descriptions of the observed events. This suggests that a discrete and structured sequence of actions serves as an effective representation for modeling video temporal dynamics.
4.4 LLM as Long-term Video Temporal Reasoner
Our experiments on LTA task reveal that compact action labels are effective for video temporal modeling, but is action representation alone sufficient for comprehensive video understanding? We then delve into a more challenging task: long-form video question answering, which demands an extensive understanding and reasoning of the video content. We conduct zero-shot experiment on EgoSchema, a recent challenging video question answering benchmark where existing state-of-the-art visual-language models struggle.
Model Input Type Full Set Acc.
InternVideo [68] frame 32.1%
GPT-4 action 38.12%
GPT-4 caption 48.26%
GPT-4* gt-narration 81.80%
Table 3: zero-shot VQA on Egoschema. * denotes subset results.
We employ OpenAI GPT-4 as the zero-shot video reasoner, and the same models used for LTA to zero-shot extract action labels and captions (BLIP-2) as well. Results are listed in Table 3. We observe the LLM utilizing both zero-shot generated actions and captions largely outperform current best-performed visual-language model (InternVideo [68]). Contrasting with the action prediction task, captions with abundant details and semantic information serve as a much stronger representation for the challenging VQA task. Furthermore, we use the video narrations of 50 examples provided by the authors as “oracle” captions and evaluate on this subset. Remarkably, the LLM achieves an impressive accuracy of 81.8% with the “oracle” captions. Although not directly comparable with the full-set performance, it reflects the vast potential and generalizability of caption-based textual video representations and the LLMs’ capability for reasoning over broad time spans.
4.5 What Makes Good Video Representation?
In addition to text features, Vamos can also integrate vision-language features. We then investigate whether different modalities encode complementary information on Ego4D, IntentQA, and Next-QA. We train the Vamos model to incorporate the visual embeddings (the linear layer in Figure 2), and use LoRA to update the weights of Llama-2.
From Ego4D LTA results shown in Table 2, we observe incorporating vision input with the compact action sequence further enhances model performance especially on verb and action prediction. For the VQA benchmarks NeXT-QA and IntentQA, as illustrated in Table 2, the caption-based textual video representations significantly outperform purely visual representations in both content description and causal reasoning. This aligns with our prior observations from EgoSchema that captions can serve as a strong representation for question answering and reasoning. Contrasting with the LTA task, adding visual features to captions only marginally influence the performance on the two VQA datasets. This suggests that the task-agnostic captions already contain sufficient information and cues to address the questions in these benchmarks.
4.6 How many tokens are needed for reasoning?
While we show zero-shot captions generated by open-box models are effective representations of video and offer a competitive interface to interact with LLMs, we further ask the following questions: How many tokens are really necessary to represent videos? What specific information from lengthy captions is actually being utilized by LLMs? We then implant a lightweight token selector to Vamos to automatically compress the captions into much more compact token sequences.
We conduct experiment on the NeXT-QA benchmark. In detail, after the word embedding layer in Vamos, the long caption-based textual video token sequence is uniformly partitioned into 40 segments. The token selector are then applied on each segment to select one token. In this way, we condense a long sequence (644 on average) into 40 tokens. These 40 tokens are then fed into LLMs as the new compressed representation of original video for down-stream video question answering task. During training, the token selector is optimized together with the Vamos.
For NeXT-QA dataset, we follow previous setting to use captions from 6 frames generated by LLaVA. For comparison, we also show the vision-only performance taking in 12 frames and the performance with unselected caption-based input. We show the results in Table 4. After condensing the prompts from an average of 644 tokens to 40 tokens, Vamos experiences an expected performance drop but still achieves a competitive 65.3% accuracy, significantly outperforming the vision-only baseline utilizing double input frames by 4.1%. When further compressed to 20 tokens, the performance drops to 59.5%, slightly worse than the vision baseline. This shows the sparsity of the free-form natural language captions for video reasoning and indicates a promising potential to compress the textual representation for better practical application. Besides, instead of using hard selection (choosing one token with highest score) during inference, we also perform soft-selection (using weighted pooling by applying Softmax instead of hard-max to the logits). Interestingly, the soft-selection achieves similar performance with hard-selection even with access to the whole sequence, indicating there is few missing information left in the unselected tokens.
 Input selected # Cau. Tem. Des. All
Frame all 61.5 58.2 66.3 61.2
Caption all 70.9 69.3 77.7 71.5
Caption 20 / 644 59.5 56.8 64.6 59.5
Caption 40 / 644 65.4 62.4 70.5 65.3
Caption* 40 / 644 65.6 62.4 70.4 65.4
Table 4: Condensing captions with token selector on NeXT-QA. Average caption length is 644. * denotes soft-selection results.
Figure 3: Illustration of Vamos predictions and manual intervention.
Figure 4: Illustration of predictions with token selector and manual intervention. The selector picks tokens highly related to the question and intervention corrects wrong predictions without training.
4.7 Qualitative Analysis and Intervention
Compared with continuous visual features, one major advantage of the textual video representation is it’s excellent interpretability that making understanding the relationship between input and prediction direct and simple. We then make qualitative analysis of some positive and negative results from Vamos and the token selector.
Visualization of results. In Figure 3 we demonstrate two correctly-answered examples and one incorrect of Vamos from IntentQA. From the two positive examples we can see that the generated captions well describe the scene and activities happening in the video (“overlooking water” and “playing keyboards”), thus providing strong clues for LLMs to answer the question about description (“look at scenery”) and reasoning (“play guitar with man playing keytar”). However, a suboptimal caption will also cause the Vamos’ failure of reasoning. In the third negative example, the captions successfully describe the baby’s action of eating but fail to acknowledge the presence and potential actions of a woman in the corner of the scene. This omission leads to an incorrect prediction regarding the woman’s reaction.
Visualization of Token Selector. By condensing the captions consists of thousands of tokens into 40 with a token selector, it not only significantly accelerates the training and inference speed, but also provides a better view for us to understand what’s the key element for Vamos to understand and reason about the video. In Figure 4, we demonstrate two positive examples and one negative example of Vamos with a token selector from NeXT-QA. After the token selector, captions in natural language are broken into separated words that highly descriptive and related to the question. The two positive examples demonstrate Vamos’ strong ability to reason by the selected tokens providing direct or indirect clues to the question. In the challenging negative example, even highly related tokens such as “race” and “track” are selected, Vamos still tends to relate “tyres” with the more common word “car”, indicating the potential bias encoded by the LLMs from pre-training.
Test Time Intervention. In addition to good interpretability, the text-based white-box representation also provides the capability of post-hoc test-time intervention, which is pivotal for diagnosing and fixing failed predictions without retraining the reasoning model. In Figure 3 and Figure 4, we show examples of test-time intervention of the negative examples for both original Vamos and Vamos with selector. By providing more accurate and related captions or tokens, Vamos is able to correct the wrong answers.
4.8 Design Choices and Ablation Study
Caption models. We study the impact of caption models on the reasoning performance on NeXT-QA. We compare two captioning models: BLIP-2 and LLaVA-1.5-7b. We observe that captions generated by BLIP-2 are generally concise spanning less than 20 tokens long while their counterparts from LLaVA-1.5 are more detailed and descriptive with about 100 tokens on average. Results shown in Table 5 shows that captions from LLaVA-1.5 achieve better performance. Interestingly, the two captions achieve similar performance on descriptive questions but LLaVA-1.5 captions perform significantly better on casual and temporal reasoning. This indicates that reasoning from videos generally benefit more from extra details compared with describing and summarizing.
Number of Captions. We study the impact of sampled frame numbers for captioning on NeXT-QA. As shown in Table 5, we found that with more captions of frames spanning across the video the model will perform better. This result is align with our expectation as we bring information with more granularity from the videos.
Comparison of LLMs on zero-shot VQA Besides, we also compare various LLMs’ zero-shot performance based on the same caption representations on EgoSchema. As shown in Table 6, GPT-4 achieves huge improvements comparing with GPT-3.5-turbo and Llama2-Chat-7B, demonstrating the potentiality and advantage of textual representation to naturally incorporate the improvements of LLMs.
 Caption Frame # Cau. Tem. Des. All
LLaVA-1.5 1 66.4 63.0 72.0 66.3
LLaVA-1.5 3 68.5 65.2 76.4 68.8
LLaVA-1.5 6 70.9 69.3 77.7 71.5
BLIP-2 6 67.9 64.6 77.1 68.4
Table 5: Ablation on captioning model and frame numbers.
4.9 Comparison with State-of-the-art
Finally, we compare our proposed Vamos with other state-of-the-art models on the four benchmarks in Tables 678 and 9. On EgoSchema zero-shot VQA, our approach based on GPT-4 outperforms the best vision-language model by 50.3%. On NeXT-QA, Vamos achieves the best performance among baselines with similiar model size. It is slightly worse than SeViLA, which trains on additional dataset for temporal localization. On IntentQA, Vamos outperforms all baselines, with a 17.4% accuracy improvement. On Ego4d LTA, Vamos outperforms AntGPT without relying on domain-specific video encoders [47].
Model Input Type Frames
VIOLET [18] frame 19.9%
FrozenBiLM [75] frame 26.9%
mPLUG-Owl [77] frame 31.1%
InternVideo [68] frame 32.1%
Llama2-13B caption 36.73%
GPT-3.5 caption 41.24%
GPT-4 caption 48.26%
Table 6: Egoschema VQA zero-shot Performance.
 Model Cau. Tem. Des. All
ATP [5] 53.1% 50.2% 66.8% 54.3%
MIST [20] 54.6% 56.6% 66.9% 57.2%
HiTeA [76] 62.4% 58.3% 75.6% 63.1%
ATM [10] 55.3% 55.6% 65.3% 57.0%
Intern Video [68] 62.5% 58.5% 75.8% 63.2%
BLIP-2 [42] 70.1% 65.2% 80.1% 70.1%
LLaMA-VQA-7B [38] 72.7% 69.2% 75.8% 72.0%
Vamos (ours) 72.6% 69.6% 78.0% 72.5%
SeViLA[79]* 74.2% 69.4% 81.3% 73.8%
Table 7: Comparison on NeXT-QA benchmark. * denotes model using additional dataset during training.
 Model CW CH TP&TN ALL
HGA [34] 44.88% 50.97% 39.62% 44.61%
HQGA [72] 48.24% 54.32% 41.71% 47.66%
VGT [73] 51.44% 55.99% 47.62% 51.27%
BlindGPT [55] 52.16% 61.28% 43.43% 51.55%
CaVIR [43] 58.4% 65.46% 50.48% 57.64%
Vamos (ours) 69.52% 70.19% 64.95% 68.51%
Table 8: Comparison with SOTA on IntentQA
 Model verb noun action
Slowfast [16] 0.717 0.736 0.925
VideoLLM [8] 0.721 0.725 0.921
PaMsEgoAI [29] 0.684 0.679 0.893
Palm [28] 0.696 0.651 0.886
AntGPT [84] 0.650 0.650 0.877
Vamos (ours) 0.643 0.650 0.868
Table 9: Comparison with SOTA on Ego4D v2 test set
5 Conclusion
We study the form of representation for long videos and proposed versatile action models (Vamos) for video understanding. Through extensive experiments on long-term action anticipation and video question answer benchmarks, we have demonstrated that surprisingly, direct application of free-form text descriptions generated by vision-language models serve as strong video representation for all benchmarks we consider, and the use of large language models enable us to perform zero-shot reasoning, and correct the answers via causal intervention. Vamos achieves state-of-the-art results on Ego4D LTA, IntentQA, close to SOTA performance on NeXT-QA, and outperforms the best vision-language model by over 50% on EgoSchema. Although our results show the promise of free-form text representation, we believe visual information is still essential for complex video understanding and reasoning, and that further explorations on modeling and benchmarking are needed.
Acknowledgements: This work is supported by Honda Research Institute USA and Samsung Advanced Institute of Technology. We would like to thank Apoorv Khandelwal, Calvin Luo, David Isele, Songpo Li, and Tian Yun for their generous feedback on this work. Our research was conducted using computational resources at Brown CCV.